OpenStudy (anonymous):

6 + 2i what is the multiplicative inverse

OpenStudy (photon336):

the multiplicative inverse means if I multiply two numbers I get 1 because \[x*\frac{ 1 }{ x } = \frac{ x }{ x } = 1\]

OpenStudy (photon336):

so the first step would be to find the multiplicative inverse that's easy it's just \[\frac{ 1 }{ 6+2i }\] now since this is a complex number we must multiply the top and bottom by the conjugate and then simplify: \[(\frac{ 1 }{ 6+2i })*(\frac{ 6-2i }{ 6-2i })\]

OpenStudy (photon336):

I'll stop here since this is sufficient to answer the question.

OpenStudy (anonymous):

thank you